admission to the senior class variety show was $3.75 for students and $4.50 for adults. The senior class collected $1680 from the sale of 400 tickets. How many student tickets were sold?
Answer by nerdybill(7384)

admission to the senior class variety show was $3.75 for students and $4.50 for adults. The senior class collected $1680 from the sale of 400 tickets. How many student tickets were sold?
.
Let x = number of student tickets sold
then
400-x = number of adult tickets sold
.
3.75x + 4.50(400-x) = 1680
3.75x + 1800 - 4.50x = 1680
1800 - 0.75x = 1680
-0.75x = -120
x = -120/(-0.75)
x = 160 (student tickets sold)